Q#1 An electron in a circular orbit about a proton can be described by classical mechanics if its angular momentum L is very much greater than h. Show that this condition is satisfied if the radius of the orbit r is very much greater than the Bohr radius a0 , i.e if

r>>a0 = 4πϵh2/e2me
1
Expert's answer
2020-06-22T11:11:55-0400

Let us consider the circular motion of electron, so the acceleration is

"a =\dfrac{v^2}{r}." The force due to which the motion exists is "F = \dfrac{1}{4\pi\varepsilon_0}\cdot\dfrac{e^2}{r^2}." Therefore, "v^2 = \dfrac{e^2}{4\pi\varepsilon_0m_er}."

The angular momentum is "L=m_evr ."

If "L \gg h," or "L^2\gg h^2, \;\;\; m_e^2v^2r^2 \gg h^2, \;\;\; m_e^2\cdot \dfrac{e^2}{4\pi\varepsilon_0m_er}\cdot r^2 \gg h^2, \;\;\; r \gg \dfrac{h^24\pi\varepsilon_0}{m_ee^2}" .